### v1

@@ 4.1 Assumption 1(i)
The continuity part of Assumption 1(i) is mild and usually trivial to verify. Under continuity, a high-level sufficient condition for existence of a maximizer of $\mathcal{G}$ over $\mathbb{R}^{d}$ is that
$\mathbb{P}\left[\limsup_{\|\mathbf{s}\|\rightarrow\infty}\mathcal{G}(\mathbf{s})<0\right]=1$ (8)
where we use $\mathbb{P}[\mathcal{G}(\mathbf{0})=0]=1$. In turn, proceeding as in the proof of Kim and Pollard (1990, Lemma 2.5) it can be shown that if the covariance kernel satisfies the (self-similarity) property that for some $H>0$,
@@
The conditions (9) and (10) are both fairly primitive. Also, by inspection (9) can be seen to hold with $H=1/2$ in each of the examples of Section 2. Moreover, setting $H=1/2$, (10) can be seen to hold with $\epsilon=3/2$ in the maximum score and empirical risk minimization examples, and with $\epsilon=1/2$ in the threshold regression example.
@@ 4.2 Assumption 1(ii)
By inspection, the displayed part of Assumption 1(ii) holds in each of the examples of Section 2. To explain why this is no coincidence, suppose the following “local uniform” version of the characterization (11) is valid: for any $\eta_{n}>0$ with $\eta_{n}=O(r_{n}^{-1})$ and any $\bm{\theta}_{n}=\bm{\theta}_{0}+O(\eta_{n})$,
Then validity of the displayed part of Assumption 1(ii) follows from the fact that
@@
$=\mathbb{E}[\{m_{n}(\mathbf{z},\bm{\theta}_{n}+\eta_{n}\mathbf{s})-m_{n}(\mathbf{z},\bm{\theta}_{n})\}\{m_{n}(\mathbf{z},\bm{\theta}_{n}+\eta_{n}\mathbf{t})-m_{n}(\mathbf{z},\bm{\theta}_{n})\}]$
for any $\mathbf{s},\mathbf{t},\mathbf{h}\in\mathbb{R}^{d}$, where $\bm{\theta}_{n}=\bm{\theta}_{0}+\eta_{n}\mathbf{h}$. Again, a characterization of the form (12) is valid in each of the examples of Section 2.
@@ 4.3.3 Example: Empirical Risk Minimization (Continued)
[…]
@@
where $u_{N}$ is a Lebesgue density of the uniform distribution on $[-N,N]^{d}$.
@@
The special structure (5) of $\mathcal{C}_{N}$ implies that $h_{N}\in\mathscr{H}_{N}$ if and only if it is of the form
$h_{N}(\mathbf{s})=\sum_{\ell=1}^{d}h_{\ell,N}(s_{\ell}),$
with $h_{\ell,N}$ belonging to the RKHS of $\mathcal{C}_{\ell,N}$ for each $\ell$. Now, proceeding as in Giné and Nickl (2016, Example 2.6.7) it can be shown that the RKHSs of $\mathcal{C}_{1,N},\dots,\mathcal{C}_{d,N}$ all (coincide and) consist of those functions on $[-N,N]$ that are zero-preserving and absolutely continuous with a square integrable (weak) derivative. In particular, we see immediately that $\mu_{N}\in\mathscr{H}_{N}$ when $\mu$ is given by (4).
@@ 5 Discussion of Assumption 1(iii)
Interpreted as a condition on the mean function $\mu$, the strength of Assumption 1 (iii) is inversely related to the richness of the RKHSs $\mathscr{H}_{N}$ generated by the covariance kernel $\mathcal{C}$. It seems natural to ask, therefore, whether certain covariance kernels are so simple that Theorem 1 is silent about the continuity properties of $F_{\hat{\mathbf{s}}}$ for many (possibly most) interesting mean functions. Revisiting a particularly simple, yet important, covariance kernel, Section 5.1 provides an affirmative answer to that question.
A related, but arguably more interesting, question is whether Assumption 1 (iii) is likely to be “close” to minimal in cases where the covariance kernel generates RKHSs that are sufficiently rich to contain many (possibly most) interesting mean functions. Revisiting another important covariance kernel, Section 5.2 provides evidence suggesting that the answer to that question will be affirmative in certain important special cases.
@@ 5.2 Two-Sided Brownian Motion
The examples in Section 2 all have the common feature that if $d=1$, then $\mathcal{C}$ is proportional to $\mathcal{C}_{\mathtt{BM}}$. More generally, regardless of the value of $d$ the examples all have the feature that $\mathcal{C}$ is a (linear) functional of $\mathcal{C}_{\mathtt{BM}}$, a feature which in turn would appear to be shared by most other examples of estimators satisfying (1) with a covariance kernel that is not bilinear. It is therefore of interest to further investigate the continuity properties of $F_{\hat{\mathbf{s}}}$ in the special case where $\mathcal{G}^{\mu}$ is a two-sided Brownian motion.

### v2

@@ 4.1 Assumption 1(i)
The continuity part of Assumption 1(i) is mild and usually trivial to verify. Under continuity and assuming that $\mathbb{P}[\mathcal{G}(\mathbf{0})=0]=1$, a high-level sufficient condition for existence of a maximizer of $\mathcal{G}$ over $\mathbb{R}^{d}$ is that
$\mathbb{P}\left[\limsup_{\|\mathbf{s}\|\rightarrow\infty}\mathcal{G}(\mathbf{s})<0\right]=1.$ (8)
In turn, proceeding as in the proof of Kim and Pollard (1990, Lemma 2.5) it can be shown that if the covariance kernel satisfies the (self-similarity) property that for some $H>0$,
@@
The assumption $\mathbb{P}[\mathcal{G}(\mathbf{0})=0]=1$ holds if (and only if) $\mu(\mathbf{0})=0=\mathcal{C}(\mathbf{0},\mathbf{0})$ and is therefore satisfied in each of the examples of Section 2. Likewise, the conditions (9) and (10) are both fairly primitive. Also, by inspection, (9) can be seen to hold with $H=1/2$ in each of the examples of Section 2. Moreover, setting $H=1/2$, (10) can be seen to hold with $\epsilon=3/2$ in the maximum score and empirical risk minimization examples, and with $\epsilon=1/2$ in the threshold regression example.
@@ 4.2 Assumption 1(ii)
By inspection, the displayed part of Assumption 1(ii) holds in each of the examples of Section 2. To explain why this is no coincidence, observe that upon defining $\bm{\theta}_{n}=\bm{\theta}_{0}+\eta_{n}\mathbf{h}$ we have
@@
$=\mathbb{E}[\{m_{n}(\mathbf{z},\bm{\theta}_{n}+\eta_{n}\mathbf{s})-m_{n}(\mathbf{z},\bm{\theta}_{n})\}\{m_{n}(\mathbf{z},\bm{\theta}_{n}+\eta_{n}\mathbf{t})-m_{n}(\mathbf{z},\bm{\theta}_{n})\}]\qquad\text{for any }\mathbf{s},\mathbf{t},\mathbf{h}\in\mathbb{R}^{d}.$
The displayed part of Assumption 1(ii) is therefore valid whenever the following “local uniform” version of (11) is valid: for any $\eta_{n}>0$ with $\eta_{n}=O(r_{n}^{-1})$ and any $\bm{\theta}_{n}=\bm{\theta}_{0}+O(\eta_{n})$,
In turn, a characterization of the form (12) is valid in each of the examples of Section 2.
@@ 4.3.3 Example: Empirical Risk Minimization (Continued)
$e_{N}(\bm{\omega};\mathbf{s})=\sqrt{24Nu_{N}(\bm{\omega})}\sum_{\ell=1}^{d}\left(\omega_{\ell}-\frac{\sqrt[3]{s_{\ell}}}{2}\right)\left[\mathbbm{1}\{0\leq\omega_{\ell}\leq\sqrt[3]{s_{\ell}}\}+\mathbbm{1}\{\sqrt[3]{s_{\ell}}\leq\omega_{\ell}<0\}\right]\sqrt{f(\theta_{0,\ell})}$
@@
where $u_{N}$ is a Lebesgue density of the uniform distribution on $N,N]^{d}$.
@@
The special structure (5) of $\mathcal{C}_{N}$ implies that $\mathscr{H}_{N}$ consists of those functions that are of the form $h_{N}(\mathbf{s})=\sum_{\ell=1}^{d}h_{\ell,N}(s_{\ell})$, with $h_{\ell,N}$ belonging to the RKHS of $\mathcal{C}_{\ell,N}$ for each $\ell$. Now, proceeding as in Giné and Nickl (2016, Example 2.6.7) it can be shown that the RKHSs of $\mathcal{C}_{1,N},\dots,\mathcal{C}_{d,N}$ all (coincide and) consist of those functions on [N,N][-N,N] that are zero-preserving and absolutely continuous with a square integrable (weak) derivative. In particular, $\mu_{N}\in\mathscr{H}_{N}$ when $\mu$ is given by (4).
@@ 5 Discussion of Assumption 1(iii)
Interpreted as a condition on the mean function $\mu$, the strength of Assumption 1 (iii) is inversely related to the richness of the RKHSs $\mathscr{H}_{N}$ generated by the covariance kernel $\mathcal{C}$. It seems natural to ask, therefore, whether certain covariance kernels are so simple that Theorem 1 is silent about the continuity properties of $F_{\hat{\mathbf{s}}}$ for many (possibly most) interesting mean functions. Revisiting a particularly simple covariance kernel, Section 5.1 provides an affirmative answer to that question.
A related, but arguably more interesting, question is whether Assumption 1 (iii) is likely to be “close” to minimal in cases where the covariance kernel generates RKHSs that are sufficiently rich to contain many (possibly most) interesting mean functions. Revisiting another well-known covariance kernel, Section 5.2 provides evidence suggesting that the answer to that question will be affirmative in certain special cases.
@@ 5.2 Two-Sided Brownian Motion
Our motivating examples have the common feature that if $d=1$, then $\mathcal{C}$ is proportional to $\mathcal{C}_{\mathtt{BM}}$. More generally, the examples have the feature that for any $d$, $\mathcal{C}$ is a (linear) functional of $\mathcal{C}_{\mathtt{BM}}$, a feature which in turn would appear to be shared by most other examples of estimators satisfying (1) with a covariance kernel that is not bilinear. It is therefore of interest to further investigate the continuity properties of $F_{\hat{\mathbf{s}}}$ in the special case where $\mathcal{G}^{\mu}$ is a two-sided Brownian motion.